Lloyd Corporation's annual pay, $1000 par value bonds, which mature in 12 years, are callable in 7 years from today at $1020, the annual coupon payment is $120. They sell at a price of $1400.

a) What is the bond's Yield to Maturity?

b) What is the bond's Current Yield?

c) What is the bond's Yield to Call?

d) What is the bond's Coupon Rate?

e) What is the bond's Selling Price if the payments are semiannual?

right answers will get a thumb's up thanks :)
Answer #1

a, PMT =120
Number of Periods =12
Par value =1000
PV =1400
YTM using financial calculator
N=12;PMT =120;PV=-1400;FV=1000;CPT I/Y =6.97%

b. Current yield =Annual Coupon/Price =120/1400 =8.57%

c. Call Price =1020
Number of years =7
YTC using financial Calculator
N=7;PMT =120;PV=-1400;FV=1040;CPT I/Y =5.46%

d. Coupon rate=Coupon/Par Value =120/1000 =12%

e. If Semi annual coupon then
PMT =120/2 =60
Number of Periods =12*2 =24
Price of Bond =PV of Coupons +PV of Par Value =60*((1-(1+6.97%/2)^-24)/(6.97%/2)+1000/(1+6.97%/2)^24 =1404.35